OpenStudy (anonymous):

Prove that 2^n+1 is not divisible by 31 posting this for the third time

OpenStudy (anonymous):

@ganeshie8

hartnn (hartnn):

i am closing this as this is feedback section. but ganeshie can continue his explanation here

ganeshie8 (ganeshie8):

lets move to math..

ganeshie8 (ganeshie8):

oh lets continue here... then :)

OpenStudy (anonymous):

oh sorry my bad

OpenStudy (anonymous):

you want me to move this there?

ganeshie8 (ganeshie8):

\(\large 2^n \equiv (2^5)^{\frac{n}{5}}\)

ganeshie8 (ganeshie8):

we oly need to consider the case n = 5k (we can get back to this later)

ganeshie8 (ganeshie8):

32 leaves a remainder 1 when divided by 31

ganeshie8 (ganeshie8):

\(\large 2^n \equiv (2^5)^{\frac{n}{5}} \equiv 1 \mod 31\)

ganeshie8 (ganeshie8):

that means, 2^n + 1 leaves a remainder 2 when divided by 31

ganeshie8 (ganeshie8):

QED

OpenStudy (anonymous):

32 leaves a remainder 1 when divided by 31 isn't this equal to 32mod31=1

ganeshie8 (ganeshie8):

yess

OpenStudy (anonymous):

how come 2^n=1mod31=1?

ganeshie8 (ganeshie8):

we're considering n = 5k case

ganeshie8 (ganeshie8):

2^n = (32)^k = (1)^k = 1 mod 31

ganeshie8 (ganeshie8):

other cases are : n=5k+1, 5k+2, 5k+3, 5k+4 they just leave a non-zero multiple of remainder of n=5k case

ganeshie8 (ganeshie8):

let me knw if congruences dont look familiar... we can avoid the congruences and try to discuss in simple math...

OpenStudy (anonymous):

yeah, can't we just avoid congruences?

ganeshie8 (ganeshie8):

we can avoid it completely if we are okay on one thing : 1) if \(a\) leaves a remainder of \(r\) when divided by \(b\), then : \(a^n\) leaves a remainder of \(r^n\) when divided by \(b\)

ganeshie8 (ganeshie8):

that observation/fact is important for the proof. i dont see any way we can avoid using it...

OpenStudy (anonymous):

is that always true? for example 5/3 will give a remainder of 2 5^2/3 will give a remainder of 1 Am I wrong?

ganeshie8 (ganeshie8):

it is always true

ganeshie8 (ganeshie8):

let me adjust ur example a bit : 5/3 will give a remainder of 2 5^2/3 will give a remainder of 2^2

ganeshie8 (ganeshie8):

\(\large 2^2 \equiv 1 \mod 3\)

ganeshie8 (ganeshie8):

when we talk about remainders when divided by 3, 1 = 4 = 7 = 10 ...

OpenStudy (anonymous):

okay now I got the whole thing.

OpenStudy (anonymous):

Really appreciate you help. Thank you for everything

ganeshie8 (ganeshie8):

np :)